# GAN decoder on a quantum toric code for noise-robust quantum teleportation

### v2

@@ Abstract
Generative adversarial network (GAN) is a strong deep learning model that has shown its value in applications such as image processing and data enhancement. Here, we propose a GAN-based quantum topological toric code decoder and we apply it to devise a quantum teleportation protocol which is robust to depolarizing noisy environments. We construct the generator and discriminator networks of GAN, train the network using the eigenvalue dataset of the toric code, and obtain an optimized decoder with a high decoding threshold compared to some existing decoders. The simulation experiments at code distances $d=3$ and $d=5$ show that the fidelity threshold of this GAN decoder is about $P=0.2108$, which is significantly larger than the threshold $P=0.1099$ of the classical decoding model. Also, the quantum teleportation protocol, optimized for noise resistance under $d=3$ and $d=5$ topological code, shows a fidelity improvement within the depolarizing noise threshold range of $P<0.06503$ and $P<0.07512$, respectively. With appropriate dataset training, the decoder can be adapted to other error models. More broadly, the proposed GAN model provides a novel approach for topological code decoders, offering a versatile framework for different types of noise processing.
@@ I Introduction
Figure 1: The basic structure of toric codes. X stabilizers are used to detect Z errors, while Z stabilizers are used to detect X errors. The measurement circuits are shown on the right. H gates, I gates, and CNOT gates are used. The ancilla qubits are measured to obtain the eigenvalue. If the data qubits have no errors, the eigenvalue will be +1; otherwise, it will be -1.
Quantum communication is an important research direction in the field of quantum information science. It realizes the interaction of information under the rules of quantum communication protocols. Quantum teleportation [23, 24, 25, 26, 27, 28] is a typical protocol in quantum communication, which has made new progress in both theory and experiment [29, 30, 31, 32]. In recent years, the achievable communication distance of quantum teleportation has been continuously upgraded. For example, in 2012, in quantum teleportation experiments using optical fibers and ground-based free-space channels, the transmission distance could reach about 100 kilometers [33]. In 2017, based on satellite platforms and space-based connections, the communication distance of quantum teleportation was extended to 1,400 kilometers [34]. In 2022, Hermans et al. used three optical connection points based on solid-state qubits to realize long-distance quantum teleportation between non-adjacent nodes in a quantum network [35]. However, due to the fragility of quantum states, noise-robust quantum teleportation methods need to be further studied [36, 37, 38, 39, 40]. Here, we address this problem.
In particular, we propose a quantum toric code with a GAN decoder and apply it to quantum teleportation. This work aims to develop a GAN decoders with higher logical fidelity compared to some existing decoders. Furthermore, the toric codes with GAN decoder are applied to quantum teleportation protocols, resulting in noise-robust and high-fidelity quantum teleportation. In this GAN decoder, two neural networks are used as generators and discriminators respectively. This decoder is trained based on the optimal decoding data set of $d=3$ and $d=5$. The trained generator generates error correction strategies with high success rates. This decoder has a higher threshold than MWPM. Specifically, under the same code distance, the error correction success rate of toric codes is higher; under the same noise rate, the decoding effect of toric codes with larger code distances is better. Here we apply the proposed GAN decoder on a quantum toric code to the teleportation protocol, but it can be extended to other quantum information protocols.
@@ II.1 Toric code
The quantum information in toric codes is distributed across multiple physical qubits on a lattice with code distance d. Larger code distances provide better fidelity within a threshold range. The core of the toric code is detecting errors by measuring operators called stabilizers. These stabilizers are divided into X stabilizers and Z stabilizers, which detect bit-flip errors and phase-flip errors, respectively. By periodically measuring these stabilizers, a set of eigenvalues, called syndromes, is obtained, enabling the detection and localization of errors.
For a toric code with a code distance of $d$, the size of the lattice $C$ is $d\times d$, and it has three main features: $f$ faces, $e$ edges, and $v$ vertices. A dual lattice ${C^{*}}$ is constructed for the toric code. The idea is to map $f$ to the dual vertices $v^{*}$, $e$ to the dual edges $e^{*}$, and $v$ to the dual faces $f^{*}$. Here $v\in C\Leftrightarrow{f^{*}}\in{C^{*}}$, $f\in C\Leftrightarrow{v^{*}}\in{C^{*}}$, $e\in C\Leftrightarrow{e^{*}}\in{C^{*}}$. The qubits of the toric code are located on $e$ and $e^{*}$. The operators at $v$ (${f^{*}}$) are the Z stabilizers, and the operators at $f$ (${v^{*}}$) are the X stabilizer, denoted as
@@
where ${\partial_{1}}e$ are the vertices of $\forall e$, ${\partial_{2}}f$ are the edges of $\forall f$. Vertices and faces anti-commute with each other, and generate the stabilizers of toric codes.
Denote the number of vertices, edges and faces of the lattice as $V$, $E$ and $F$, respectively. The toric code has $V+F-2$ independent generators. It follows from the stabilizer theory [4] that the number of encoded qubits is
@@
where $\chi$ is the Euler characteristic, $g$ is the number of closed surface loops. A toric code with a code distance of $d$ has $V=F=d^{2}$, $2d^{2}$ qubits and $22d^{2}-$ 2 stabilizers.
@@ II.3 Quantum teleportation
Figure 3: Quantum teleportation. Particles $A$ and $B$ are the two particles of the EPR pair. The state of particle $C$, denoted as $\left|\psi_{C}\right\rangle$, is to be teleported. Alice holds particles $A$ and $C$. Bob holds particle $B$. The gray rectangles represent quantum operations.
@@ III.1 Error correction
Assume Pauli errors $E_{p}={X_{c}}{Z_{{c^{*}}}}$ follow a probability distribution $p=\left\{{{p_{c,{c^{*}}}}}\right\}$. The first step in error correction is to measure the stabilizers of the toric code and obtain the syndrome. $Z_{v}$ and $X_{f}$ are used to detect the X error and Z error of the qubits, respectively. If errors occur, the eigenvalues of corresponding stabilizers change from $+1$ to $-1$, which means a pair of $-1$ defects are generated on the vertices of the lattice. There are ${d^{2}}-1$ independent operators on a $d\times d$ lattice , thus a single $-1$ eigenvalue with all other $+1$ is impossible. The probability to obtain a given syndrome is
$P\left({\partial c,{\partial^{*}}{c^{*}}}\right)=\sum\limits_{e\in{X_{c}}}{\sum\limits_{{e^{*}}\in{Z_{{c^{*}}}}}{P\left({c+e,c+{e^{*}}}\right)}}.$ (3)
The syndrome is degenerate. For error correction, we need to select a set of correction qubits to eliminate the defects in the syndrome in pairs. If the correction qubits and error qubits form a trivial loop, the logical state of the toric code will return to its original state. However, if they form a nontrivial loop around the lattice, the error correction will fail. Although the nontrivial loop eliminates defects, it alters the original state of the toric code. An optimal decoder selects the chain with the minimum amount of correction qubits. The success probability of this optimal strategy is
@@
Figure 4: Error correction of a quantum toric code. Red data qubits indicate that an error has occurred. Blue data qubits mean they are selected as an error correction qubit. (a) A toric code with error data qubits. The eigenvalues of the stabilizers around the error qubits becomes $-1$. (b) Failed quantum toric code error correction. A error correction chain and the error chain form a nontrivial loop. (c) Successful quantum toric code error correction. A error correction chain and the error chain form a trivial loop. (d) Another successful quantum toric code error correction. This strategy forms a trivial loop, but not the optimal.
For a string of errors, we need to find an error correction strategy. Since the error correction in toric codes is degenerate, there is not just one right correction scheme for a given set of error syndromes. This is different from other quantum error correction codes, where each error has a fixed correction strategy. In toric codes, error correction is successful as long as the error qubits and correction qubits form a trivial loop. Conversely, forming a non-trivial loop means error correction has failed. Thus, the goal of our decoder is to generate a correction path such that the qubits on this path and the error qubits form a trivial loop.
@@
Here we consider a depolarizing noise model, where X and Z errors occur independently with a certain probability $P$. Correcting independent X errors on the lattice C and correcting independent Z errors on the dual lattice $C^{*}$ is completely equivalent. In this GAN decoder, the focus is on generating the result of defect matching elimination, thus training a decoder model for one type of error is sufficient.
@@ III.2.2 Structure of the generator
$\text{ReLU}(x)=\begin{cases}x,&\text{if }x>0\\ 0,&\text{if }x\leq 0\end{cases},$ (7)
@@ III.2.3 Structure of the discriminator
$\text{LReLU}(x)=\begin{cases}x,&\text{if }x>0\\ \alpha x,&\text{if }x\leq 0\end{cases},$ (11)
@@ III.3.3 Error correction process
If no error occurs, due to the commutation between the stabilizers of the same kind, the eigenvalues of $Z_{v}$ and $X_{f}$ are stable, $\hat{Z}=+1$, $\hat{X}=+1$. On the contrary, if errors occur, due to the anti-commutation between errors and stabilizers, the eigenvalues of adjacent stabilizers change from $+1$ to $-1$, $\hat{Z}=-1$, $\hat{X}=-1$.
@@ IV.1 Model training
When the learning rate is set to 0.0002, the loss value decreases very slowly, requiring more time to achieve the desired result. At a learning rate of 0.02, the loss value decreases rapidly, but oscillations occur, which affect the network’s convergence. When the learning rate is 0.002, the convergence is better: the loss value decreases quickly and remains stable. Therefore, we set the learning rate to 0.002 for training.
Next, we employ Adam to train the decoder, which utilizes the first and second momentum terms during training. This makes it a robust optimization technique for non-smooth objective functions and noisy gradients [47]. According to Eqs. (14) and (15), we use the loss function to show whether our method converged. We calculate the distance between target data and generate data in the feature space using the Frechet Inception Distance (FID) [48]
@@
[…]
Figure 9 shows the training results of the GAN decoder. Figures 9a and 9b display the convergence of the generator and discriminator losses. The GAN decoder models for the quantum toric codes with $d=3$ and $d=5$ were considered sufficiently trained after 500 and 1800 training iterations, respectively. Additionally, Figures 9c and 9d illustrate the changes in FID with iterations, demonstrating how the distribution of the generated data evolves toward the distribution of the target data. After 500 and 1800 training iterations for $d=3$ and $d=5$ respectively, the FID approaches convergence, indicating that the distribution of the generated samples is close to that of the training samples. Table 4 shows the average decoding time and the average training time per iteration for GAN decoders with $d=3$ and $d=5$.
@@
Code distance Avg. decoding time Avg. training time / iteration
@@ IV.2 Result evaluation
Through training of the GAN model, we achieved a logical fidelity of 99.895% for the quantum toric code decoder. We tested the error correction capabilities of the trained GAN model and the local MWPM model on toric code lattices with code distances of $d=3$ and $d=5$. Considering that the encoding scheme of toric codes is not unique, all comparative experiments in our study use the same encoding scheme to ensure fairness and consistency. The test results are presented in Figures 10 and 11, illustrating the optimization performance and the logarithmic behavior of the GAN decoder in comparison to the MWPM decoder. As the error rate increase, the error correction success rate significantly decreases. The impact of code distance on logical fidelity depends on the threshold. The threshold of a quantum toric code decoder is a critical indicator that defines the maximum tolerable physical error rate under which logical error rates can be effectively suppressed by increasing the code distance. Below the threshold, the logical error rate decreases as the code distance grows. Conversely, above the threshold, the logical error rate either stops improving or even increases, rendering error correction ineffective.
The logical fidelity threshold for the local MWPM model is approximately $P=0.1099$, similar to the RL decoder [12], while the threshold for our GAN decoder model is approximately $P=0.2108$. Overall, in terms of logical fidelity performance, the GAN model outperforms these existing decoders for decoding toric codes with $d=3$ and $d=5$.
@@
Error rate (×$\times 10^{-3}$) Original fidelity (%) Number of failures at d=5 ($n_{\mathrm{shoot}}=1024$) Optimizing fidelity (%)
@@
Figures 13 and 14 compare the optimization of the teleportation method using topological codes with the GAN decoder and show the logarithmic behavior of the solution. The figure indicates that in the depolarizing noise simulation experiment, for the system with $d=3$ toric code, the fidelity threshold of the model is approximately 0.06503; for the system with $d=5$ toric code, the fidelity threshold is approximately 0.07512. This implies that the model can ensure effective fidelity improvement in quantum networks with error rates below the threshold in the noise model system.

### v3

@@ Abstract
We propose a generative adversarial network (GAN)-based decoder for quantum topological codes and apply it to enhance a quantum teleportation protocol under depolarizing noise. By constructing and training the GAN’s generator and discriminator networks using eigenvalue datasets from the code, we obtain a decoder with a significantly improved decoding pseudo-threshold. Simulation results show that our GAN decoder achieves a pseudo-threshold of approximately $p=0.2108$, estimated from the crossing point of logical error rate curves for code distances $d=3$ and $d=5$, nearly double that of a classical decoder under the same conditions ($p\approx 0.1099$). Moreover, at the same target logical error rate, the GAN decoder consistently achieves higher logical fidelity compared to the classical decoder. When applied to quantum teleportation, the protocol optimized using our decoder demonstrates enhanced fidelity across noise regimes. Specifically, for code distance $d=3$, fidelity improves within the depolarizing noise threshold range $P<0.06503$; for $d=5$, the range extends to $P<0.07512$. Moreover, with appropriate training, our GAN decoder can generalize to other error models. This work positions GANs as powerful tools for decoding in topological quantum error correction, offering a flexible and noise-resilient framework for fault-tolerant quantum information processing.
@@ I Introduction
Figure 1: The basic structure of toric codes. X stabilizers are used to detect Z errors, while Z stabilizers are used to detect X errors. The measurement circuits are shown on the right. H gates, I gates, and CNOT gates are used. The ancilla qubits are measured to obtain the eigenvalue. For a given stabilizer, the eigenvalue is +1 if the number of errors on the neighboring data qubits is even, and -1 if it is odd.
Quantum communication is an important research direction in the field of quantum information science. It realizes the interaction of information under the rules of quantum communication protocols. Quantum teleportation [23, 24, 25, 26, 27, 28] is a typical protocol in quantum communication, which has made new progress in both theory and experiment [29, 30, 31, 32]. In recent years, the achievable communication distance of quantum teleportation has been continuously upgraded. For example, in 2012, in quantum teleportation experiments using optical fibers and ground-based free-space channels, the transmission distance could reach about 100 kilometers [33]. In 2017, based on satellite platforms and space-based connections, the communication distance of quantum teleportation was extended to 1,400 kilometers [34]. In 2022, Hermans et al. used three optical connection points based on solid-state qubits to realize long-distance quantum teleportation between non-adjacent nodes in a quantum network [35]. However, due to the fragility of quantum states, effective quantum error correction and decoding methods are essential for improving the reliability of quantum teleportation [36, 37, 38, 39, 40, 41, 42]. Here, we investigate decoding under a toric code and apply it to quantum teleportation to enhance its robustness under depolarizing noise.
In particular, we propose a novel GAN decoder for the quantum toric code. The primary aim of this work is to develop a decoder that achieves higher logical fidelity compared to some existing decoding methods. Furthermore, we demonstrate the application of the toric code with the GAN decoder in a quantum teleportation protocol, illustrating its potential to enable more noise-robust and high-fidelity quantum teleportation. In this GAN decoder, two neural networks are used as generators and discriminators respectively. The trained generator generates error correction strategies with high success rates. This decoder has a higher pseudo-threshold and fidelity than MWPM. Specifically, below the threshold, toric codes have a higher error correction success rate at the same code distance, and larger code distances provide better decoding performance under the same noise rate. Here we apply the proposed GAN decoder on a quantum toric code to the teleportation protocol, but it can be extended to other quantum information protocols.
@@ II.1 Toric code
The quantum information in toric codes is distributed across multiple physical qubits on a lattice with code distance d. Larger code distances achieve better fidelity below the threshold. The core of the toric code is detecting errors by measuring operators called stabilizers. These stabilizers are divided into X stabilizers and Z stabilizers, which detect bit-flip errors and phase-flip errors, respectively. By periodically measuring these stabilizers, a set of eigenvalues, called syndromes, is obtained, enabling the detection and localization of errors.
For a toric code with a code distance of $d$, the size of the lattice $C$ is $d\times d$, and it has three main features: $f$ faces, $e$ edges, and $v$ vertices. A dual lattice $C^{*}$} is constructed for the toric code. The idea is to map $f$ to the dual vertices $v^{*}$, $e$ to the dual edges $e^{*}$, and $v$ to the dual faces $f^{*}$. Here $v\in C\Leftrightarrow{f^{*}}\in{C^{*}}$, $f\in C\Leftrightarrow{v^{*}}\in{C^{*}}$, $e\in C\Leftrightarrow{e^{*}}\in{C^{*}}$. The data qubits of the toric code are located on $e$ and $e^{*}$. The operators at $v$ ($f^{*}$}) are the Z stabilizers, and the operators at $f$ ($v^{*}$}) are the X stabilizer, denoted as
@@
where ${\partial_{1}}e$ are the vertices of $\forall e$, ${\partial_{2}}f$ are the edges of $\forall f$. Vertices and faces anti-commute with each other, and generate the stabilizers of toric codes. This anti-commutation relation ensures that bit-flip and phase-flip errors are independently detectable and correctable. It also guarantees that the stabilizer group forms a valid quantum code space, which encodes logical qubits in topologically protected entangled states.
Denote the number of vertices, edges and faces of the lattice as $V$, $E$ and $F$, respectively. The toric code has V+F2V+F-2 independent generators. It follows from the stabilizer theory [4] that the number of encoded qubits is
@@
where $\chi$ is the Euler characteristic, $g$ is the number of closed surface loops. A toric code with a code distance of $d$ has $V=F=d^{2}$, $2d^{2}$ qubits and $2d^{2}-2$ stabilizers.
@@ II.3 Quantum teleportation
Figure 3: Quantum teleportation. Particles AA and $B$ are the two particles of the EPR pair. The state of particle $C$, denoted as $\left|\psi_{C}\right\rangle$, is to be teleported. Alice holds particles $A$ and $C$. Bob holds particle $B$. The gray rectangles represent quantum operations.
@@ III.1 Error correction
Assume $E_{p}={X_{c}}{Z_{{c^{*}}}}$ represents a Pauli error, where $c$ denotes a chain (subset) of edges on the primal lattice indicating bit-flip errors, and $c^{*}$ denotes a chain of edges on the dual lattice indicating phase-flip errors. $X_{c}$ is the product of Pauli-X operators acting on the edges in $c$, and $Z_{c^{*}}$ is the product of Pauli-Z operators acting on the edges in $c^{*}$. The joint probability distribution is given by $p=\left\{{{p_{c,{c^{*}}}}}\right\}$. The first step in error correction is to measure the stabilizers of the toric code and obtain the syndrome. $Z_{v}$ and $X_{f}$ are used to detect the X error and Z error of the qubits, respectively. If errors occur, the eigenvalues of corresponding stabilizers change from +1+1 to 1-1, which means a pair of 1-1 defects are generated on the vertices of the lattice. There are ${d^{2}}-1$ independent operators on a $d\times d$ lattice , thus a single 1-1 eigenvalue with all other +1+1 is impossible. The probability to obtain a given syndrome is
$P\left({\partial c,{\partial^{*}}{c^{*}}}\right)=\sum\limits_{e\in{X_{c}}}{\sum\limits_{{e^{*}}\in{Z_{{c^{*}}}}}{P\left({c+e,{c^{*}}+{e^{*}}}\right)}}.$ (3)
The syndrome is degenerate. For error correction, we need to select a set of correction qubits to eliminate the defects in the syndrome in pairs. If the correction qubits and error qubits form a trivial loop, the logical state of the toric code will return to its original state. However, if they form a nontrivial loop around the lattice, the error correction will fail. Although the nontrivial loop eliminates defects, it alters the original state of the toric code. The MWPM decoder selects the chain with the minimum amount of correction qubits. The success probability of this strategy is
@@
Figure 4: Error correction of a quantum toric code. Red data qubits indicate that an error has occurred. Blue data qubits mean they are selected as an error correction qubit. (a) A toric code with error data qubits. The eigenvalues of the stabilizers around the error qubits becomes 1-1. (b) Failed quantum toric code error correction. A error correction chain and the error chain form a nontrivial loop. (c) Successful quantum toric code error correction. A error correction chain and the error chain form a trivial loop. (d) Another successful quantum toric code error correction. This strategy forms a trivial loop, but not the optimal.
For a string of errors, we need to find an error correction strategy. Since the error correction in toric codes is degenerate, there is not just one right correction scheme for a given set of error syndromes. This differs from classical non-degenerate error correction codes, where each error typically has a unique correction strategy. In toric codes, error correction is successful as long as the error qubits and correction qubits form a trivial loop. Conversely, forming a non-trivial loop means error correction has failed. Thus, the goal of our decoder is to generate a correction path such that the qubits on this path and the error qubits form a trivial loop.
@@
Here we consider a depolarizing noise model, in which an input qubit undergoes a completely random Pauli error ($X$, $Y$, or $Z$) with a certain probability $p$, and with probability 1p1-p the qubit remains unchanged. Mathematically, it can be written as
${\cal E}\left(\rho\right)=\left({1-p}\right)\rho+\frac{p}{3}\left({X\rho X+Y\rho Y+Z\rho Z}\right),$ (5)
where $\rho$ is the input qubit state, $p$ is the depolarizing probability, and $X$, $Y$, $Z$ are the Pauli operators.
Correcting independent X errors on the lattice C and correcting independent Z errors on the dual lattice $C^{*}$ is completely equivalent. In this GAN decoder, the focus is on generating the result of defect matching elimination, thus training a decoder model for one type of error is sufficient.
@@ III.2.2 Structure of the generator
[…]
@@ III.2.3 Structure of the discriminator
[…]
@@ III.3.3 Error correction process
If no error occurs, due to the commutation between the stabilizers of the same kind, the eigenvalues of $Z_{v}$ and $X_{f}$ are stable, $\hat{Z}=+1$, $\hat{X}=+1$. On the contrary, if errors occur, due to the anti-commutation between errors and stabilizers, the eigenvalues of adjacent stabilizers change from +1+1 to 1-1, $\hat{Z}=-1$, $\hat{X}=-1$.
@@ IV.1 Model training
Figure 9: Training performance at different learning rates. The black dotted line shows the loss function at a learning rate of 0.0002; the red line at 0.002; and the blue line at 0.02.
@@
Figure 9 shows the training performance at different learning rates. When the learning rate is set to 0.0002, the loss value decreases very slowly, requiring more time to achieve the desired result. At a learning rate of 0.02, the loss value decreases rapidly, but oscillations occur, which affect the network’s convergence. When the learning rate is 0.002, the convergence is better: the loss value decreases quickly and remains stable. Therefore, we set the learning rate to 0.002 for training.
Next, we employ the Adam optimization algorithm to train the network parameters. Adam uses both the first and second moment estimates during training, making it a robust optimization technique for non-smooth objective functions and noisy gradients [49]. According to Eqs. (15) and (16), we use the loss function to show whether our method converged. We calculate the distance between target data and generate data in the feature space using the Frechet Inception Distance (FID) [50]
@@
where $\mu_{t}$} and $\mu_{g}$} are the averages of the feature vectors of the target data and the generated data, respectively. $C_{t}$} and $C_{g}$} are their respective covariance matrices. $\mathrm{Tr}\left[\cdot\right]$ denotes the trace of a matrix. $\left\|\cdot\right\|$ denotes the $L_{2}$ norm of a vector.
Figure 10 shows the training results of the GAN decoder. Figures 10a and 10b display the convergence of the generator and discriminator losses. The GAN decoder models for the quantum toric codes with $d=3$ and $d=5$ were considered sufficiently trained after 500 and 1800 training iterations, respectively. After averaging over samples, the resulting loss values typically fall within the range $(0,2)$. In our training, the final loss value converges to approximately 0.7, which reflects a balanced optimization between the generator and discriminator networks. This value indicates that the model has learned to generate outputs that are reasonably close to the target distribution while maintaining sufficient discriminator sensitivity. Additionally, Figures 10c and 10d illustrate the changes in FID with iterations, demonstrating how the distribution of the generated data evolves toward the distribution of the target data. After 500 and 1800 training iterations for $d=3$ and $d=5$ respectively, the FID approaches convergence, indicating that the distribution of the generated samples is close to that of the training samples.
Specifically, the generator of the GAN decoder contains approximately 409,153 trainable parameters, while the discriminator includes about 2,914,433 parameters. Table 4 shows the average decoding time per sample and the average training time per iteration for GAN decoders with code distances d=3 and d=5. As the code distance increases, both the training time and the decoding time increase. Although the number of trainable parameters remains fixed, larger code distances result in higher computational costs and greater memory usage.
As the code distance increases, the size of the input syndrome and the GAN networks grows, leading to higher computational and memory requirements. In practice, the training time scales roughly quadratically to cubically with code distance, depending on the network architecture and dataset size. However, once trained, the model can be reused for multiple decoding tasks, and incremental fine-tuning can reduce retraining costs for slightly larger distances.
During inference, the GAN decoder performs a single forward pass through the generator network for each syndrome. Therefore, the decoding time scales approximately linearly with the number of qubits (proportional to the square of the code distance for surface codes).
@@
Code distance Avg. decoding time / sample Avg. training time / iteration
@@ IV.2 Result evaluation
By training the GAN model, we achieved a logical fidelity of 99.895% at $p=0.05$ for the toric code decoder with code distance $d=5$. We tested the error correction capabilities of the trained GAN model and the local MWPM model on toric code lattices with code distances $d=3$ and $d=5$. Considering that the encoding scheme of toric codes is not unique, all comparative experiments in our study use the same encoding scheme to ensure fairness and consistency. The test results are presented in Figures 11 and 12, illustrating the optimization performance and the logarithmic behavior of the GAN decoder in comparison to the MWPM decoder. As the error rate increase, the error correction success rate significantly decreases.
The impact of code distance on logical fidelity can be characterized by the pseudo-threshold. In this work, the pseudo-threshold of the quantum toric code decoder is estimated based on the intersection point of the logical error rate curves for code distances d=3 and d=5. This crossing point provides only an approximate finite-size estimate for the depolarizing noise model and higher than the theoretical threshold of the toric code, which is known to be below 11%[44]. Below the threshold, the logical error rate decreases as the code distance grows. Conversely, above the threshold, the logical error rate either stops improving or even increases, rendering error correction ineffective.
Under the same conditions, the pseudo-threshold for the local MWPM model is approximately $P=0.1099$, which is close to that of the RL decoder [12] and the RNN decoder ($P=0.1363$) [11], while our GAN-based decoder achieves a significantly higher pseudo-threshold of $P=0.2108$. Overall, in terms of logical fidelity performance, the GAN model outperforms these existing decoders for decoding toric codes with $d=3$ and $d=5$.
@@
Error rate ($\times 10^{-3}$) Original fidelity (%) Number of failures at d=5 ($n_{\mathrm{shoot}}=1024$) Optimizing fidelity (%)
@@
Figures 14 and 15 compare the optimization of the teleportation method using topological codes with the GAN decoder and show the logarithmic behavior of the solution. The figure indicates that in the depolarizing noise simulation experiment, for the system with $d=3$ toric code, the fidelity threshold of the model is approximately 0.06503; for the system with $d=5$ toric code, the fidelity threshold is approximately 0.07512. This implies that the model can ensure effective fidelity improvement in quantum networks with error rates below the threshold in the noise model system.